# Can Large Language Models Act as Ensembler for Multi-GNNs?

### v1

@@ Appendix A Baselines
(2) Graph Neural Networks: GCN (Kipf and Welling, 2016) is a fundamental method based on convolutional neural networks which operates directly on graph-structured data. GAT (Veličković et al., 2017) computes the hidden representations of each node in the graph by first learning the importance of its neighbors and then aggregating information from them. GIN (Xu et al., 2019) develop a simple architecture that is provably the most expressive among the class of GNNs and is as powerful as the Weisfeiler-Lehman graph isomorphism test. GraphSAGE (Hamilton et al., 2017a) present a general inductive framework that leverages node feature information (e.g., text attributes) to efficiently generate node embeddings for previously unseen data. Graphormer (Ying et al., 2021) is built upon the standard Transformer (Vaswani, 2017) architecture, and could attain excellent results on a broad range of graph representation learning tasks. Graphormer propose several simple yet effective structural encoding methods to help Graphormer better model graph-structured data. GT (Dwivedi and Bresson, 2012) presented a simple yet effective approach to generalize transformer networks on arbitrary graphs. It includes an elegant way to fuse node positional features using Laplacian eigenvectors for graph datasets, inspired from the heavy usage of positional encodings in NLP transformer models and recent research on node positional features in GNNs. CoarFormer (Kuang et al., 2024) is a two-view architecture that captures fine-grained local information using a GNN-based module on the original graph and coarse yet long-range information using a Transformer-based module on the coarse graph.
@@
(3) Language Models: BERT (Devlin, 2018) is a groundbreaking model in natural language processing. It utilizes the transformer architecture to understand the context of words in a sentence by looking at both their left and right contexts simultaneously. This bidirectional approach enables BERT to capture nuanced meanings and relationships in text. SentenceBERT (Reimers, 2019) is a modification of the original BERT (Bidirectional Encoder Representations from Transformers) model designed specifically for generating sentence embeddings. It was introduced to improve the performance of BERT on tasks that require understanding the semantic meaning of entire sentences, such as semantic textual similarity, paraphrase identification, and clustering. DeBERTa (He et al., 2020) is a transformer-based language model that improves upon BERT and other models like RoBERTa by introducing several innovations.
@@
(4) LLM Based: DGTL (Qin et al., 2024) incorporates graph structure information through tailored disentangled graph neural network layers, enabling LLMs to capture the intricate relationships hidden in TAGs from multiple structural factors, which is able to enhance the reasoning and predicting capabilities of LLMs for TAGs. SNS-GPT4 (Li et al., 2024) is a specialized version of the GPT-4 model designed for Social Network Services (SNS) applications. It introduces similarity-based neighbor selection to effectively improve the quality of selected neighbors, thereby enhancing graph representation and alleviating issues like over-squashing and heterophily. InstructGraph-INS (Wang et al., 2024) is an advanced framework that integrates instruction-based learning with graph neural networks (GNNs) to improve performance in various graph-based tasks. The model leverages both explicit instructions and implicit graph structures to enhance representation learning and decision-making processes. InstructGLM (Ye et al., 2024) use natural language to describe multi-scale geometric structure of the graph and then instruction finetune an LLM to perform graph tasks, which enables Generative Graph Learning. GAugLLM (Fang et al., 2024) is a novel framework for augmenting TAGs. It leverages advanced large language models like Mistral to enhance self-supervised graph learning. Baichuan2 (Yang et al., 2023) refers to the second iteration of the Baichuan large language model (LLM). It has been trained on an impressive 2.6 trillion high-quality tokens, ensuring a robust understanding of language nuances. Baichuan 2 comes in two main versions: 7B and 13B, both available in Base and Chat configurations, with the latter offering 4-bit quantization for efficient deployment.
Appendix B GNN Encoders
Graph Convolutional Networks (GCNs) GCNs extend the concept of convolutional neural networks (CNNs) to graph data. They operate by aggregating and transforming feature information from a node’s local neighborhood, effectively capturing the graph’s structural information. The propagation rule in GCNs typically involves normalizing the adjacency matrix to ensure proper scaling of node features. GCNs have been widely used in tasks such as node classification, link prediction, and graph classification due to their ability to learn representations that incorporate both node features and graph topology.
Graph Attention Networks (GATs) GATs introduce an attention mechanism to the graph convolutional framework, allowing the model to learn the relative importance of neighboring nodes when aggregating their features. This is achieved by computing attention coefficients for each pair of connected nodes and using these coefficients to perform a weighted aggregation of the neighbor features. The attention mechanism in GATs makes them more flexible and capable of handling graphs with varying degrees of connectivity and node importance. GATs have shown strong performance in various graph-based tasks, providing a way to dynamically adjust the focus on relevant neighbors.
Graph Isomorphism Networks (GINs) GINs are designed to address the limitations of previous GNN models in distinguishing graph structures. They aim to be as powerful as the Weisfeiler-Lehman graph isomorphism test, which is a method for determining whether two graphs are isomorphic. GINs use a simple yet effective aggregation function that sums the features of a node’s neighbors and applies a MLP to the result. This approach ensures that GINs can capture complex graph structures and provide highly discriminative node and graph embeddings. GINs are particularly effective in graph classification tasks, where capturing subtle differences in graph topology is crucial.
Appendix C Detailed hyperparameters
The detailed hyperparameters of the first and second stages of LensGNN  are shown in Tables 6 and 7, respectively.
Table 6. Detailed hyperparameters of aligning multi-GNNs.
Model GCN GAT GIN
Dataset Cora Pubmed ogb-arxiv Cora Pubmed ogb-arxiv Cora Pubmed ogb-arxiv
Hidden size 256 256 256 256 256 256 256 256 256
GNN Representation size 40960 40960 20480 40960 40960 20480 40960 40960 20480
Number of layers 2 2 2 2 2 2 2 2 2
Learning rate of GNN encoder 0.0002 0.0002 0.01 0.0002 0.0002 0.01 0.0001 0.0001 0.004
Initial learning rate of the classifier 0.0002 0.0002 0.0005 0.0002 0.0002 0.0005 0.0001 0.0001 0.0005
Dropout Rate 0.25 0.25 0.2 0.25 0.25 0.2 0.25 0.25 0.2
Weight Decay 0.25 0.25 0 0.25 0.25 0 0.25 0.25 0
@@ Appendix C Detailed hyperparameters
Table 7. Detailed hyperparameters of Ensembling multi-GNNs with LLM (LoRA fine-tune).
Model Baichuan2-13B Falcon-7B InternLM2.5-7B-chat
Dataset Cora Pubmed ogb-arxiv Cora Pubmed Cora Pubmed
Lora dropout 0.1 0.1 0.1 0.1 0.1 0.1 0.1
Loraplus Ir ratio 16 16 16 16 16 16 16
Prompt cutoff length 2047 2047 2047 2047 2047 2047 2047
Training batch size per device 4 4 4 4 4 4 4
Gradient accumulation steps 1 1 1 1 1 1 1
Learning rate 0.0001 0.00005 0.0001 0.0001 0.0001 0.0001 0.0001
Early stop epoch 3 2 3 3 2 3 2
Warmup Ratio 0.1 0.1 0.1 0.1 0.1 0.1 0.1
Parameter precision FP16 FP16 FP16 FP16 FP16 FP16 FP16

### v2

@@ Appendix B Baselines
(2) Graph Neural Networks: GCN Kipf and Welling (2016) is a fundamental method based on convolutional neural networks which operates directly on graph-structured data. GAT Veličković et al. (2017) computes the hidden representations of each node in the graph by first learning the importance of its neighbors and then aggregating information from them. GIN Xu et al. (2019) develop a simple architecture that is provably the most expressive among the class of GNNs and is as powerful as the Weisfeiler-Lehman graph isomorphism test. GraphSAGE Hamilton et al. (2017a) present a general inductive framework that leverages node feature information (e.g., text attributes) to efficiently generate node embeddings for previously unseen data. Graphormer Ying et al. (2021) is built upon the standard Transformer Vaswani (2017) architecture, and could attain excellent results on a broad range of graph representation learning tasks. Graphormer propose several simple yet effective structural encoding methods to help Graphormer better model graph-structured data.
@@
(3) LM Based Models: BERT Devlin (2018) is a groundbreaking model in natural language processing. It utilizes the transformer architecture to understand the context of words in a sentence by looking at both their left and right contexts simultaneously. This bidirectional approach enables BERT to capture nuanced meanings and relationships in text. SentenceBERT Reimers (2019) is a modification of the original BERT (Bidirectional Encoder Representations from Transformers) model designed specifically for generating sentence embeddings. It was introduced to improve the performance of BERT on tasks that require understanding the semantic meaning of entire sentences, such as semantic textual similarity, paraphrase identification, and clustering. DeBERTa He et al. (2020) is a transformer-based language model that improves upon BERT and other models like RoBERTa by introducing several innovations.
Figure 5: A prompt template for node classification task.
Figure 6: A prompt template for graph classification task.
@@
(4) LLM Based Models: GraphGPT Tang et al. (2024) is a novel framework that integrates Large Language Models (LLMs) with graph structural knowledge through graph instruction tuning. This innovative approach enables LLMs to comprehend and interpret the structural components of graphs, thereby demonstrating superior generalization in both supervised and zero-shot graph learning tasks. LLaGA Chen et al. (2024) integrates the capabilities of LLMs with graph-structured data, enabling LLMs to handle complex graph tasks effectively. LLaGA achieves this by reorganizing graph nodes into structure-aware sequences and mapping them into the token embedding space through a versatile projector, demonstrating superior versatility, generalizability, and interpretability across various datasets and tasks. InstructGLM Ye et al. (2024) use natural language to describe multi-scale geometric structure of the graph and then instruction finetune an LLM to perform graph tasks, which enables Generative Graph Learning. DGTL Qin et al. (2024) incorporates graph structure information through tailored disentangled graph neural network layers, enabling LLMs to capture the intricate relationships hidden in TAGs from multiple structural factors, which is able to enhance the reasoning and predicting capabilities of LLMs for TAGs. SNS-GPT4 Li et al. (2024) is a specialized version of the GPT-4 model designed for Social Network Services (SNS) applications. It introduces similarity-based neighbor selection to effectively improve the quality of selected neighbors, thereby enhancing graph representation and alleviating issues like over-squashing and heterophily. GAugLLM Fang et al. (2024) is a novel framework for augmenting TAGs. It leverages advanced large language models like Mistral to enhance self-supervised graph learning. Baichuan2 Yang et al. (2023) refers to the second iteration of the Baichuan large language model (LLM). It has been trained on an impressive 2.6 trillion high-quality tokens, ensuring a robust understanding of language nuances. Baichuan 2 comes in two main versions: 7B and 13B, both available in Base and Chat configurations, with the latter offering 4-bit quantization for efficient deployment. OFA Liu et al. (2024) is a pioneering framework that unifies various graph tasks into a single model, enabling it to address classification tasks across different domains and tasks. OFA achieves this by converting graph data into text-attributed graphs (TAGs), using language models to encode diverse text attributes into a common embedding space, and introducing a novel graph prompting paradigm for in-context learning without fine-tuning. GOFA Kong et al. (2024) is a novel graph foundation model that integrates the strengths of large language models (LLMs) and graph neural networks (GNNs) to enable joint graph and language modeling. It achieves this by interleaving GNN layers into a frozen pre-trained LLM, allowing for self-supervised pretraining on graph data and fluidity in handling various graph-related tasks.
Table 8: Hyperparameter Search Spaces.
Hyperparameter Search space
learning rate $\{2.0\times 10^{-5},5.0\times 10^{-5},1.0\times 10^{-4}\}$
dropout $[0,0.25]$
loraplus lr ratio $\{16.0,24.0,32.0\}$
Graph token $t$$\{1,2,4,8,16\}$
Table 9: Detailed hyperparameters of aligning multi-GNNs.
Model Dataset Hidden size GNN representation size Number of layers Learning rate of GNN encoder Initial learning rate of the classifier Dropout Rate Weight Decay
GCN Cora 256 40960 2 0.0002 0.0002 0.25 0.001
Pubmed 256 40960 2 0.0002 0.0002 0.25 0.001
ogbn-arxiv 256 20480 2 0.01 0.0005 0.2 0
Citeseer 256 40960 2 0.0001 0.0005 0.25 0.001
Wiki-CS 256 40960 2 0.0002 0.0005 0.25 0.001
BACE 256 40960 4 0.0001 0.0005 0.1 0.0005
BBBP 256 40960 2 0.0001 0.0005 0.1 0.0005
ClinTox 256 40960 2 0.0001 0.0005 0.1 0.0005
GAT Cora 256 40960 2 0.0002 0.0002 0.25 0.001
Pubmed 256 40960 2 0.0002 0.0002 0.25 0.001
ogbn-arxiv 256 20480 2 0.01 0.0005 0.2 0
Citeseer 256 40960 2 0.0001 0.0005 0.25 0.001
Wiki-CS 256 40960 2 0.0002 0.0005 0.25 0.001
BACE 256 40960 4 0.0001 0.0005 0.1 0.0005
BBBP 256 40960 2 0.0001 0.0005 0.1 0.0005
ClinTox 256 40960 2 0.0001 0.0005 0.1 0.0005
GIN Cora 256 40960 2 0.0001 0.0001 0.25 0.001
Pubmed 256 40960 2 0.0001 0.0001 0.25 0.001
ogbn-arxiv 256 20480 2 0.004 0.0005 0.2 0
Citeseer 256 40960 2 0.00005 0.0005 0.25 0.001
Wiki-CS 256 40960 2 0.0001 0.0005 0.25 0.001
BACE 256 40960 4 0.0001 0.0005 0.1 0.0005
BBBP 256 40960 2 0.0001 0.0005 0.1 0.0005
ClinTox 256 40960 2 0.0001 0.0005 0.1 0.0005
@@
Table 10: Detailed hyperparameters of Ensembling multi-GNNs with LLM (LoRA fine-tune).
Model Dataset Lora dropout Loraplus lr ratio Training batch size Learning rate Early stop epoch Warmup ratio
Baichuan2-13B Cora 0.1 16 4 0.0001 3 0.1
Pubmed 0.1 16 4 0.00005 2 0.1
ogbn-arxiv 0.1 16 4 0.0001 3 0.1
Citeseer 0.15 16 4 0.0001 5 0.1
Wiki-CS 0.1 16 4 0.0001 5 0.1
BACE 0.1 16 4 0.0001 4 0.1
BBBP 0.1 16 4 0.0001 3 0.1
ClinTox 0.1 16 4 0.0001 3 0.1
Baichuan2-7B Cora 0.15 16 4 0.0001 3 0.1
Pubmed 0.15 16 4 0.0001 2 0.1
ogbn-arxiv 0.15 16 4 0.0001 3 0.1
InternLM2.5-7B Cora 0.1 16 4 0.0001 3 0.1
Pubmed 0.1 16 4 0.0001 2 0.1
ogbn-arxiv 0.1 16 4 0.0001 3 0.1
Falcon-7B Cora 0.1 16 4 0.0001 3 0.1
Pubmed 0.1 16 4 0.0001 2 0.1
Appendix C Experiment setup
We implement LensGNN by PyTorch. We primarily use Baichuan2 Yang et al. (2023) and InternLM2.5 Cai et al. (2024) as the backbone LLM for LensGNN. In the experiments, we ensemble three widely adopted GNN models: GCN, GAT and GIN, each with two layers. We perform grid search to fine-tune hyperparameters based on the validation set. Details on the search space is given in Table 8. We utilize LoRA+ Hayou et al. (2024) for fine-tuning. Some fixed LoRA settings include capping each training sample at 2,047 tokens and using half-precision (FP16) for LoRA fine-tuning, with a batch size of 4 per GPU and gradient updates every step. We utilize a cosine-type learning rate scheduler and set the warmup ratio to 0.1. For the training of LLMs, we utilize LLaMA-Factory Zheng et al. (2024) as a framework. For GNN models, we use node representations obtained from the pre-trained SentenceBERT Reimers (2019) as input. For baselines that report results on the adopted datasets, we directly report the results from their original papers. For those whose results are missing, we leave them blank. For Baichuan2, we directly use it without fine-tuning. We run all the experiments on a server with a single NVIDIA Tesla A100 GPU.
Appendix D Prompt design
The prompt is divided into two parts: Instruction and Input. The former specifies what the LLM should do with the input, defines the format of the input, and highlights the characteristics of the input content. The latter provides the specific text of target node or graph and their multiple GNN representations. An example of a prompt template can be seen in Figure 5. For graph classification, we follow MolXPT Liu et al. (2023) to design prompts for molecular graphs as shown in Figure 6.
Appendix E Detailed hyperparameter settings
The details on the setting of hyperparameters in the first and second stages of LensGNN  are shown in Tables 9 and 10, respectively. The hyperparameters for the 20-shot experiment are shown in Tables 11 and 12.
Table 11: Detailed hyperparameters of aligning multi-GNNs (20-shot).
Model Dataset Hidden size GNN representation size Number of layers Learning rate of GNN encoder Initial learning rate of the classifier Dropout Rate Weight Decay
GCN Cora 256 32768 2 0.0002 0.0005 0.25 0.001
Pubmed 256 32768 2 0.0002 0.0005 0.25 0.001
GAT Cora 256 32768 2 0.0002 0.0005 0.25 0.001
Pubmed 256 32768 2 0.0002 0.0005 0.25 0.001
GIN Cora 256 32768 2 0.0001 0.0005 0.25 0.001
Pubmed 256 32768 2 0.0001 0.0005 0.25 0.001
Table 12: Detailed hyperparameters of Ensembling multi-GNNs with LLM (LoRA fine-tune) (20-shot).
Model Dataset Lora dropout Loraplus lr ratio Training batch size Learning rate Early stop epoch Warmup ratio
Baichuan2-7B Cora 0.15 16 4 0.0001 8 0.1
Pubmed 0.15 16 4 0.0001 9 0.1
InternLM2.5-7B Cora 0.1 16 4 0.0001 8 0.1
Pubmed 0.05 16 4 0.0001 9 0.1